# Large-scale, Independent and Comprehensive study of the power of LLMs for test case generation

### v1

@@ Abstract.
Unit testing, crucial for identifying bugs in code modules like classes and methods, is often neglected by developers due to time constraints. Automated test generation techniques have emerged to address this, but often lack readability and require developer intervention. Large Language Models (LLMs), like GPT and Mistral, show promise in software engineering, including in test generation. However, their effectiveness remains unclear.
This study conducts the first comprehensive investigation of LLMs, evaluating the effectiveness of four LLMs and five prompt engineering techniques, for unit test generation. We analyze 216 300 tests generated by the selected advanced instruct-tuned LLMs for 690 Java classes collected from diverse datasets. We assess correctness, understandability, coverage, and bug detection capabilities of LLM-generated tests, comparing them to EvoSuite, a popular automated testing tool. While LLMs show potential, improvements in test correctness are necessary. This study reveals the strengths and limitations of LLMs compared to traditional methods, paving the way for further research on LLMs in software engineering.
@@ 1. Introduction
Unit tests plays a crucial role in software development, as they validate individual functions or code units, ensuring their correctness and reliability (Beck, 2000; Siddiqui, 2021; Shore and Warden, 2021). In addition to verification, unit tests help in early bug detection, enhance code quality, provide documentation, support debugging, and reduce the risk of costly errors in production. However, manual unit test creation can be arduous and discouraging, leading some developers to skip testing entirely (Daka and Fraser, 2014; Kochhar et al., 2013b, a).
@@
To cope with the challenges of test creation by developers, automated test generation techniques have gained traction, including methods like fuzzing (Miller et al., 1990; Zeller et al., 2019), feedback-directed random test generation (Toffola et al., 2017; Csallner and Smaragdakis, 2004; Pacheco and Ernst, 2007; Pacheco et al., 2008; Selakovic et al., 2018; Arteca et al., 2022), dynamic symbolic execution (Godefroid et al., 2005; Sen et al., 2005; Cadar et al., 2008; Tillmann et al., 2014), and search-based and evolutionary approaches (Fraser and Arcuri, 2011a, b). These techniques aim to create tests that maximize coverage through static or dynamic analyses. EvoSuite (Fraser and Arcuri, 2011b), a prominent search-based method, uses evolutionary algorithms to generate test suites for Java classes, aiming to maximize coverage. However, despite its effectiveness, the generated tests often lack readability (Almasi et al., 2017; Grano et al., 2018; Roy et al., 2020), contain obscure variable names (Daka et al., 2015), lack assertions (Panichella et al., 2020) or include overly general or excessive assertions (Panichella et al., 2020), and have shortcomings in detecting real bugs (Bacchelli et al., 2008; Pinto and Vergilio, 2010; Shamshiri et al., 2015; Almasi et al., 2017). Although these tests can inspire manual development of high-coverage test suites, they often lack naturalness and practical usability (Yuan et al., 2023; Shin et al., 2023; Almasi et al., 2017; Panichella et al., 2020; Alagarsamy et al., 2023).
@@
Given the limitations of traditional analysis-based test generation methods, recent research has focused on using advanced deep learning techniques to improve unit test generation. Some techniques treat test generation as a neural machine translation problem, fine-tuning models to translate classes or methods into corresponding test prefixes or assertions (Tufano et al., 2020; Watson et al., 2020; Dinella et al., 2022; Nie et al., 2023; Shin et al., 2023; Alagarsamy et al., 2023). Recently, Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), have been leveraged for this purpose due to their extensive parameter sets and versatility (Fan et al., 2023). Studies (Wang et al., 2024; Fan et al., 2023; Kang et al., 2023; Lemieux et al., 2023; Rao et al., 2023; Chen et al., 2023; Siddiq et al., 2023) have then shown that LLMs can significantly enhance the quality of automated test case generation, streamlining the process for developers. These LLM-based test generation approaches are claimed to address issues of readability, understandability, and meaningfulness (Siddiq et al., 2024b; Yuan et al., 2023; Tang et al., 2024). The emergence of LLMs therefore may mark a significant evolution in software development, enabling more automated, intelligent, and efficient testing processes.
@@
Evaluation methodology: Beyond classical metrics such as code coverage, few works actually attempt to assess whether the LLM-generated test cases are readable or actually contribute to bug detection.
This paper: Our work investigates the effectiveness of LLMs in unit test generation and how prompt engineering techniques might affect their performance. Drawing inspiration from (1) works that compare Search-Based Software Testing (SBST) against ChatGPT (Tang et al., 2024), (2) comprehensive surveys in the field (Wang et al., 2024), and (3) insightful early experiments on the threats of using LLMs in software engineering (Sallou et al., 2023), we undertake a large-scale study to evaluate the potential of LLMs and prompt engineering to improve unit test case generation.
Unlike much of the prior research, which focus mainly on criteria related to syntax accuracy, compilability, and at best, code coverage, our study further evaluates these criteria along with the capability of LLMs to generate functional test suites that can identify real-world software bugs.
@@
We describe our findings based on a comprehensive array of evaluation metrics, including correctness, readability, code coverage, and the capacity for bug detection. By leveraging various datasets and models, we provide a robust basis for comparisons and future research. In particular, our assessment on an “unseen” dataset strengthens our study’s relevance and mitigates biases related to data leakage in LLMs.
@@
The remainder of this paper is structured as follows: Section 2 details the background of LLMs for unit test generation, and prompt engineering. Section 3 outlines our experimental setup, followed by Section 4, which discusses the evaluation results. While Section 5 addresses potential study threats and limitations, Section 6 reviews related work, and Section 7 offers concluding remarks.
@@ 2.1. LLM-based Unit Test Generation
Large language models (LLMs) are extensive neural networks built primarily on the Transformer architecture (Vaswani et al., 2017) and trained on vast text corpora and code using self-supervised techniques like Masked Language Modeling and Causal Language Modeling (Shanahan, 2024; Naveed et al., 2023; Hadi et al., 2023; Shanahan, 2024). These models, including encoder-only like BERT (Devlin et al., 2018), decoder-only like GPT (Brown et al., 2020; Yenduri et al., 2024), and encoder-decoder models like T5 (Raffel et al., 2020), are often enhanced with methods like Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Bai et al., 2022) to better align with human intentions. When interacting with these models, the instructions provided alongside any input data form what is referred to as a prompt (Siddiq et al., 2024b; Amatriain, 2024; Vogelsang and Fischbach, 2024; Siddiq et al., 2024a). Indeed, a prompt is a text description in natural language that specifies the task to be performed by an LLM. As shown in Figure 1, an LLM can be prompted 222https://chatgpt.com/share/99d6468e-98ed-40b1-9b14-17bb9c239f9c to generate a test suite for a Divider Class, and it will respond by producing the test suite, illustrating its ability to understand and execute the task specified in the prompt.
@@ 2.2. Prompt Engineering
Prompt engineering involves creating prompts which enable LLMs to effectively address specific tasks (Amatriain, 2024; Siddiq et al., 2024a; Vogelsang and Fischbach, 2024; Naveed et al., 2023; Sahoo et al., 2024). This technique is particularly useful in complex problem-solving scenarios, where a single line of reasoning might not be enough (Amatriain, 2024; Kojima et al., 2022; Wei et al., 2022; Reynolds and McDonell, 2021). It allows LLMs to adopt a more human-like approach by considering multiple possibilities before reaching a conclusion. This emerging field involves crafting prompts that guide LLMs to generate desired outputs, ranging from broad descriptions to detailed specifications. The importance of prompt engineering is twofold: Firstly, highly specific prompts yield precise responses, while more general prompts may produce less relevant outputs (Si et al., 2022). Secondly, adjusting prompt instructions can help control the output format of LLMs, minimizing the need for post-processing and improving performance evaluation (Macedo et al., 2024).
@@
Various studies (Fan et al., 2023; Wang et al., 2024) have leveraged LLMs and prompt engineering for automated software engineering tasks, commonly using basic approaches such as Zero-shot (Kojima et al., 2022) and Few-shot learning (Brown et al., 2020). Recently, research has broadened prompt engineering to include techniques such as Chain-of-Thought (Wei et al., 2022), self-consistency (Wang et al., 2022), and Tree-of-Thoughts (Wang et al., 2022). Figure 2 illustrates various examples of prompts (additional examples of prompt engineering techniques are available in the replication package).
@@
Zero-shot learning enables models to generalize to new tasks without specific examples but may falter with complex tasks (Sahoo et al., 2024; Kojima et al., 2022). Few-shot learning, or in-context learning, uses a few examples to improve model performance on novel tasks (Brown et al., 2020; Reynolds and McDonell, 2021). Chain-of-Thought (CoT), introduced by Wei et al. (Wei et al., 2022), outlines intermediate reasoning steps and can be enhanced by combining it with few-shot prompts for complex reasoning. Self-consistency, proposed by Wang et al. (Wang et al., 2022), uses diverse reasoning paths to select the most consistent answers, improving deep reasoning tasks. Tree-of-Thoughts (ToT), a newer technique proposed by Yao et al. (Wang et al., 2022) and (Long, 2023), builds on Chain-of-Thought by employing a systematic exploration through a ”tree” of reasoning steps, integrating search algorithms for more strategic problem-solving. This study explores various prompt engineering methods to adapt pre-trained LLMs for unit test generation, emphasizing their application in guiding LLM behavior in test generation tasks.
@@ 3.1. Overview
Figure 3 outlines our methodology, detailing the steps from prompting LLMs using various datasets to generate unit test cases, through to evaluating these tests. The process involves generating test suites for each class, repeated thirty times. Syntactically correct test suites are isolated, compiled, and validated, discarding any incorrect or textual outputs. The first two research questions focus on analyzing LLM-generated tests in terms of output format, syntax correction, compilability (RQ1), and comprehensibility (RQ2). The last two questions compare the performance of LLM-generated test suites with those created by EvoSuite in terms of code coverage (RQ3) and bug detection (RQ4).
@@ 3.2. Research Questions
RQ4: How effective are LLM-generated test suites in detecting bugs compared to EvoSuite-generated test suites? This research question examines the practical value of LLM-generated test suites. We compare their ability to uncover software bugs against the ability of EvoSuite, a well-established SBST technique. High code coverage, which is a known strength of SBST techniques, is crucial, but it doesn’t guarantee efficient bug detection. Therefore, with this research, we aim to identify if LLMs offer any strengths in finding software bugs beyond the capabilities of traditional SBST techniques.
@@ 3.3. Large Language Models
One such model, GPT-3.5-turbo, has a token limit of 4,096. Research suggests that exceeding this limit when splitting a class into test case generation methods can hinder the model’s understanding of the class context, potentially reducing test quality (Tang et al., 2024). To ensure a fair comparison, all models in our study are restricted to this token limit, and any inputs exceeding it are excluded. Token counts are determined using OpenAI’s tiktoken333https://github.com/openai/tiktoken library.
@@
| Custom mini Dataset (CMD) | Includes ODC a collaborative project to enhance database functionalities and Conductor OSS an open-source project management tool aimed at efficiency and scalability. Combines bugged and fixed classes for bug detection analysis. | 2 | 31 | 20 | 16 | 9 | 5 |
@@ 3.4. Prompt Design
Unit test generation research highlights the importance of practical and focused methods that avoid overly intricate solutions (e.g., (Siddiq et al., 2024b; Yuan et al., 2023; Chen et al., 2023; Tang et al., 2024)). To achieve this, we designed prompts with varying complexity levels based on a clear and well-structured template. Each prompt consists of two parts: a Natural Language Description (NLD) outlining the task and a Source Code placeholder to insert the Class Under Test (CUT). We will now explore the details of each part, explaining how they align with established unit testing practices.
@@
This study investigates the impact of varying prompt complexity on the quality of generated test cases. We employ five distinct prompt engineering techniques: Zero-shot Learning (ZSL), Few-shot learning (FSL), Chain-of-Thought (CoT), Tree-of-thoughts (ToT), and a novel technique called Guided Tree-of-thoughts (GToT). ZSL provides prompts without specific examples, forcing models to generalize to new tasks. FSL offers a few examples to guide the model towards better performance. CoT structures prompts to encourage logical reasoning, while ToT promotes exploration of multiple reasoning paths for more informed decisions. GToT attempts to exploit advantages of both CoT and ToT by guiding the reasoning of each solicited expert within the prompt itself. This research focuses on these techniques as test case generation demands a deep understanding of the code and the ability to reason through complex scenarios. Further details and comparisons of these prompt templates are available in the replication package.

### v2

@@ Abstract.
Unit testing, crucial for ensuring the reliability of code modules, such as classes and methods, is often overlooked by developers due to time constraints. Automated test generation techniques have emerged to address this, but they frequently lack readability and require significant developer intervention. Large Language Models (LLMs), such as GPT and Mistral, have shown promise in software engineering tasks, including test generation, but their overall effectiveness remains unclear.
This study presents an extensive investigation of LLMs, evaluating the effectiveness of four models and five prompt engineering techniques for unit test generation. We analyze 216 300 tests generated by the selected advanced instruct-tuned LLMs for 690 Java classes collected from diverse datasets. Our evaluation considers correctness, understandability, coverage, and test smell detection in the generated tests, comparing them to a widely used automated testing tool, EvoSuite. While LLMs demonstrate potential, improvements in test quality—particularly in reducing common test smells—are necessary. This study highlights the strengths and limitations of LLM-generated tests compared to traditional methods, paving the way for further research on LLMs in test automation.
@@ 1. Introduction
Unit tests plays a crucial role in software development, as they validate individual functions or code units, ensuring their correctness and reliability (Beck, 2000; Siddiqui, 2021; Shore and Warden, 2021). In particular, unit tests help in early bug detection, enhance code quality, provide documentation, support debugging, and reduce the risk of costly errors in production. However, manual unit test creation can be arduous and discouraging, leading some developers to skip testing entirely (Daka and Fraser, 2014; Kochhar et al., 2013b, a).
@@
To cope with the challenges of test creation by developers, automated test generation techniques have gained traction, including methods like fuzzing (Miller et al., 1990; Zeller et al., 2019), feedback-directed random test generation (Toffola et al., 2017; Pacheco and Ernst, 2007; Pacheco et al., 2008), dynamic symbolic execution (Godefroid et al., 2005; Sen et al., 2005; Cadar et al., 2008), and search-based and evolutionary approaches (Fraser and Arcuri, 2011a, b). These techniques aim to create tests that maximize coverage through static or dynamic analyses. EvoSuite (Fraser and Arcuri, 2011b), a prominent search-based method, uses evolutionary algorithms to generate test suites for Java classes, aiming to maximize coverage. However, despite its effectiveness, the generated tests often lack readability (Almasi et al., 2017; Grano et al., 2018; Roy et al., 2020), contain obscure variable names (Daka et al., 2015), lack assertions (Panichella et al., 2020) or include overly general or excessive assertions (Panichella et al., 2020), and have shortcomings in detecting real bugs (Bacchelli et al., 2008; Pinto and Vergilio, 2010; Shamshiri et al., 2015; Almasi et al., 2017). Although these tests can inspire manual development of high-coverage test suites, they often lack naturalness and practical usability (Yuan et al., 2023; Shin et al., 2023; Almasi et al., 2017; Panichella et al., 2020; Alagarsamy et al., 2023).
@@
Given the limitations of traditional analysis-based test generation methods, recent research has focused on using advanced deep learning techniques to improve unit test generation. Some techniques treat test generation as a neural machine translation problem, fine-tuning models to translate classes or methods into corresponding test prefixes or assertions (Tufano et al., 2020; Watson et al., 2020; Dinella et al., 2022; Nie et al., 2023; Shin et al., 2023; Alagarsamy et al., 2023). Recently, Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), have been leveraged for this purpose due to their extensive parameter sets and versatility (Fan et al., 2023). Studies (Wang et al., 2024; Fan et al., 2023; Kang et al., 2023; Lemieux et al., 2023; Chen et al., 2023; Siddiq et al., 2023) have then shown that LLMs can significantly enhance the quality of automated test case generation, streamlining the process for developers. These LLM-based test generation approaches are claimed to address issues of readability, understandability, and meaningfulness (Siddiq et al., 2024b; Yuan et al., 2023; Tang et al., 2024). The emergence of LLMs therefore may mark a significant evolution in software development, enabling more automated, intelligent, and efficient testing processes.
@@
Evaluation methodology: Beyond classical metrics such as code coverage, few works actually attempt to assess whether the LLM-generated test cases are readable or contribute to the presence of test smells. In our study, we go beyond these traditional metrics by focusing on the prevalence of test smells, which are indicative of poor design and maintainability in test suites.
This paper: Our work investigates the effectiveness of LLMs in unit test generation and how prompt engineering techniques might affect their performance. Drawing inspiration from (1) works that compare Search-Based Software Testing (SBST) against ChatGPT (Tang et al., 2024; Siddiq et al., 2024b), (2) comprehensive surveys in the field (Wang et al., 2024), and (3) insightful early experiments on the threats of using LLMs in software engineering (Sallou et al., 2023), we undertake a large-scale study to evaluate the potential of LLMs and prompt engineering to improve unit test case generation.
Unlike much of the prior research, which focuses mainly on criteria related to syntax accuracy, compilability, and at best, code coverage, our study further evaluates these criteria along with the capability of LLMs to generate test suites that are free from common test smells.
@@
We describe our findings based on a comprehensive array of evaluation metrics, including correctness, readability, code coverage, and the prevalence of test smells. By leveraging various datasets and models, we provide a robust basis for comparisons and future research. In particular, our assessment on an ”unseen” dataset strengthens our study’s relevance and mitigates biases related to data leakage in LLMs.
@@
The remainder of this paper is structured as follows: Section 2 details the background of LLMs for unit test generation, and prompt engineering. Section 3 outlines our experimental setup, followed by Section 4, which discusses the evaluation results. While Section 5 addresses potential study threats and limitations, Section 6 reviews related work. Section 7 discusses the findings, their implications, and takeaways. Section 8 offers concluding remarks.
@@ 2.1. LLM-based Unit Test Generation
Large language models (LLMs) are extensive neural networks built primarily on the Transformer architecture (Vaswani et al., 2017) and trained on vast text corpora and code using self-supervised techniques like Masked Language Modeling and Causal Language Modeling (Shanahan, 2024; Naveed et al., 2023; Hadi et al., 2023). These models, including encoder-only like BERT (Devlin et al., 2018), decoder-only like GPT (Brown et al., 2020; Yenduri et al., 2024), and encoder-decoder models like T5 (Raffel et al., 2020), are often enhanced with methods like Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Bai et al., 2022) to better align with human intentions. When interacting with these models, the instructions provided alongside any input data form what is referred to as a prompt (Siddiq et al., 2024b; Amatriain, 2024; Vogelsang and Fischbach, 2024; Siddiq et al., 2024a). Indeed, a prompt is a text description in natural language that specifies the task to be performed by an LLM. As shown in Figure 1, an LLM can be prompted 222https://chatgpt.com/share/99d6468e-98ed-40b1-9b14-17bb9c239f9c to generate a test suite for a Divider Class, and it will respond by producing the test suite, illustrating its ability to understand and execute the task specified in the prompt.
@@ 2.2. Prompt Engineering
Prompt engineering involves creating prompts which enable LLMs to effectively address specific tasks (Amatriain, 2024; Siddiq et al., 2024a; Vogelsang and Fischbach, 2024; Naveed et al., 2023; Sahoo et al., 2024). This technique is particularly useful in complex problem-solving scenarios, where a single line of reasoning might not be enough (Amatriain, 2024; Kojima et al., 2022; Wei et al., 2022). It allows LLMs to adopt a more human-like approach by considering multiple possibilities before reaching a conclusion. This emerging field involves crafting prompts that guide LLMs to generate desired outputs, ranging from broad descriptions to detailed specifications. The importance of prompt engineering is twofold: Firstly, highly specific prompts yield precise responses, while more general prompts may produce less relevant outputs (Si et al., 2022). Secondly, adjusting prompt instructions can help control the output format of LLMs, minimizing the need for post-processing and improving performance evaluation (Macedo et al., 2024).
@@
Various studies (Fan et al., 2023; Wang et al., 2024) have leveraged LLMs and prompt engineering for automated software engineering tasks, commonly using basic approaches such as Zero-shot (Kojima et al., 2022) and Few-shot learning (Brown et al., 2020). Recently, research has broadened prompt engineering to include techniques such as Chain-of-Thought (Wei et al., 2022), self-consistency (Wang et al., 2022), and Tree-of-Thoughts (Wang et al., 2022). Figure 2 provides examples of the prompts used for test generation, which were carefully crafted following prompt engineering best practices and insights from prior research (Tang et al., 2024; Chen et al., 2023; Tufano et al., 2020) (additional examples of prompt engineering techniques are available in the replication package).
@@
Zero-shot learning enables models to generalize to new tasks without specific examples but may falter with complex tasks (Sahoo et al., 2024; Kojima et al., 2022). Few-shot learning, or in-context learning, uses a few examples to improve model performance on novel tasks (Brown et al., 2020; Reynolds and McDonell, 2021). Chain-of-Thought (CoT), introduced by Wei et al. (Wei et al., 2022), outlines intermediate reasoning steps and can be enhanced by combining it with few-shot prompts for complex reasoning. Self-consistency, proposed by Wang et al. (Wang et al., 2022), uses diverse reasoning paths to select the most consistent answers, improving deep reasoning tasks. Tree-of-Thoughts (ToT), a newer technique proposed by Yao et al. (Yao et al., 2024) and (Long, 2023), builds on Chain-of-Thought by employing a systematic exploration through a ”tree” of reasoning steps, integrating search algorithms for more strategic problem-solving. This study explores various prompt engineering methods to adapt pre-trained LLMs for unit test generation, emphasizing their application in guiding LLM behavior in test generation tasks.
2.3. Test Smells
The primary goal of unit testing is to ensure that individual components of a program function as expected and meet specified requirements (Koomen and Pol, 1999). A unit typically refers to an isolated segment of code, such as a method or class, that can be tested independently. Like production code, unit tests need to be not only correct but also maintainable and readable (Gonzalez et al., 2017). Test smells are indicators of potential issues or poor practices within a unit test suite (Greiler et al., 2013; Guerra and Fernandes, 2007; Palomba et al., 2016; Peruma et al., 2019). These smells may signal inefficiencies, complexity, or excessive coupling to the underlying code, making the tests harder to maintain or understand (Meszaros et al., 2003). Examples include tests that are overly fragile, difficult to understand, or tightly bound to implementation details. For instance, the Java code in Listing 1 and Listing 2 from the TsDetect documentation 333https://testsmells.org/pages/testsmellexamples.html highlights two common test smells: Assertion Roulette and Magic Number Test. These test smells frequently appear in automatically generated tests, such as those produced by tools like EvoSuite or LLM-based test generation.
Assertion Roulette occurs when multiple assertions are made in a test without clear explanations, making it difficult to identify which assertion caused the test to fail. In Listing 1, the method assertThat() is called three times, each testing a different condition, but without any explanation messages. If one of these assertions fails, identifying the source of the failure becomes challenging. Magic Number Test refers to the use of unexplained numeric literals in test assertions, making it unclear what the values represent. In Listing 2, the test uses numeric literals in the assertions (15 and 30) without explaining their significance. While the test may be syntactically correct and compilable, the lack of clarity hinders maintainability and understandability.
Although these tests are functionally correct, their test smells make them harder to maintain and debug, presenting challenges in long-term software development.
@@ 3.1. Overview
Our methodology, illustrated in Figure 3, involves prompting LLMs with diverse datasets to generate unit test cases, followed by evaluation for syntax correctness, compilability, and validation. Any substandard outputs are discarded, and the process is repeated thirty times to ensure reliability. To ensure a fair comparison, we adopted the same approach as EvoSuite, which was executed thirty times for each class in (Tang et al., 2024).
Our study focuses on generating unit tests at the class level, reflecting real-world testing scenarios where testing an entire class provides a more comprehensive evaluation of the model’s capabilities. Class-level testing captures the interactions between methods within a class and presents more complex challenges than method-level testing. By focusing on the class level, we aim to assess how well LLMs can generate tests that account for dependencies and interactions within a class. We investigate the quality of LLM-generated tests in terms of output format, syntax correctness, and compilability (RQ1), as well as comprehensibility (RQ2). We also compare them with EvoSuite in terms of code coverage (RQ3) and test smell detection (RQ4).
@@ 3.2. Research Questions
RQ4: How do LLM-generated test suites compare to EvoSuite-generated test suites in terms of test smell prevalence? This research question explores the quality of LLM-generated test suites by analyzing the presence of test smells—indicators of poor design and maintainability. We compare the absence of test smells of LLM-generated test suites against those produced by EvoSuite. The goal is to determine whether LLMs, through prompt engineering, can produce tests that are less prone to common test smells and thus more maintainable than those generated by traditional SBST techniques like EvoSuite.
@@ 3.3. Large Language Models
One such model, GPT-3.5-turbo, has a token limit of 4,096. Research suggests that exceeding this limit when splitting a class into test case generation methods can hinder the model’s understanding of the class context, potentially reducing test quality (Tang et al., 2024). To ensure a fair comparison, all models in our study are restricted to this token limit, and any inputs exceeding it are excluded. Token counts are determined using OpenAI’s tiktoken444https://github.com/openai/tiktoken library.
@@
| Custom mini Dataset (CMD) | Includes ODC a collaborative project to enhance database functionalities and Conductor OSS an open-source project management tool aimed at efficiency and scalability. | 2 | 31 | 20 | 16 | 9 | 5 |
@@ 3.4. Prompt Design
Unit test generation research highlights the importance of practical and focused methods that avoid overly complicated solutions (e.g., (Siddiq et al., 2024b; Yuan et al., 2023; Chen et al., 2023; Tang et al., 2024)). To achieve this, we designed prompts with varying complexity levels based on a clear and well-structured template. Each prompt consists of two parts: a Natural Language Description (NLD) outlining the task and a Source Code placeholder to insert the Class Under Test (CUT). We will now explore the details of each part, explaining how they align with established unit testing practices.
@@
This study investigates the impact of varying prompt complexity on the quality of generated test cases. We employ five distinct prompt engineering techniques: Zero-shot Learning (ZSL), Few-shot learning (FSL), Chain-of-Thought (CoT), Tree-of-thoughts (ToT), and a novel technique called Guided Tree-of-thoughts (GToT). ZSL provides prompts without specific examples, forcing models to generalize to new tasks. FSL offers a few examples to guide the model towards better performance. For Few-shot learning, we chose a medium-sized human-written class and its test suite to fit within the LLM token limit. This pair was used consistently across all experiments. CoT structures prompts to encourage logical reasoning, while ToT promotes exploration of multiple reasoning paths for more informed decisions. GToT attempts to exploit advantages of both CoT and ToT by guiding the reasoning of each solicited expert within the prompt itself. The main intuition is that combining the strengths of CoT and ToT in a single prompt could improve LLMs performances. While CoT instructs a single tester to list methods, generate basic and edge case tests, and compile them into a JUnit file, GToT simulates three testers collaborating to refine test cases at each step, resulting in more comprehensive and accurate tests. This study focuses on these techniques as test case generation demands a deep understanding of the code and the ability to reason through complex scenarios. Further details and comparisons of these prompt templates are available in the replication package.